# Review of Demographic Fairness in Face Recognition

### v2

@@ IV Assessment of Demographic Fairness in FR
Since the 2019 Face Recognition Vendor Test (FRVT) [^2] Since 2023, the FRVT initiative has been restructured into the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) programs. report, NIST has included demographic effects in FR algorithms [15]. This assessment involved comparing non-mated pairs within the same demographic group, setting thresholds for algorithms to achieve an FMR of 0.0001 for white males (since this demographic typically associated with the lowest FMR). The report [15] offered a comprehensive analysis of recognition processes and identified areas where demographic effects might occur. To quantify demographic disparity, NIST initially employed Inequity Ratios (IR), calculating the ratio of maximum to minimum of FMR and FNMR, at score threshold $\tau$, across demographic groups as shown in Eq. 1.
@@
–where, the subscripts $d_{i},d_{j}$ refer to the demographic group; and $\alpha,\beta$ are weighing coefficients.
However, considering potentially large range of the error rates, these ratios can become numerically unstable, especially in extreme cases. To alleviate this shortcoming, NIST has considered few modified versions of Inequity Ratios such as adjusting the score threshold ($\tau$), incorporating fixed constants in the denominator, or expressing worst-case error rates relative to arithmetic or geometric means [85]. Using the geometric mean is particularly advantageous due to its extended range over FMR/FNMR values. Another possible approach involves referencing a standard FMR or FNMR in the denominator, which inherently resolves stability issues while providing more robust evaluations of demographic fairness.
@@
Schuckers et al. highlighted the importance of accounting for statistical variation when evaluating fairness in FR systems [87]. They noted that the differences among demographic groups can arise either from actual performance disparities or by chance due to sampling variability, leading to potential Type-I errors. To address this, they proposed two statistical methodologies: a bootstrap-based hypothesis test and a simpler test methodology tailored for non-statistical audiences. Their study also conducted simulations to explore the relationship between margin of error and factors such as the number of subjects, attempts, correlation between attempts, underlying FNMRs, and the number of demographic groups. In [88], the authors propose a statistical framework for assessing fairness using Receiver Operating Characteristics (ROC) curves and associated uncertainty measures derived from U-statistics. They argue that the ROC curve is not only central to evaluating system performance but also serves as a key tool for analyzing fairness in similarity scoring. They established asymptotic guarantees for empirical ROC curves and fairness-relevant metrics, emphasizing the importance of accurately quantifying uncertain. Notably, they highlight that conventional bootstrap methods may misrepresent uncertainty due to the U-statistical nature of FMR and FNMR, and recommend a recentering technique to ensure valid inference. In [90], the researchers at MdTF proposed new metric for assessment of bias based on demographic outcomes. Their metric, GARBE (Gini Aggregation Rate for Biometric Equitability), is inspired by the Gini coefficient– which has a long history of use as a dispersion measure in socio-economic context. The GARBE evaluates statistical dispersion in error rates and emphasizes equitable treatment across demographic groups. Similar to the FDR and IR, this metric combines weighted contributions of FMR and FNMR to produce a single fairness score for a given fixed score threshold. The formula for GABRE and its constituent Gini coefficient are provided in Eq. 3.
@@
The Sum of Group Error Differences (SEDG) was introduced as the fairness assessment metric address disparities in biometric verification systems in [91]. The SEDG calculates relative deviations in FMR and FNMR across demographic groups from the FMR/ FNMR of global scores. They consider the Equal Error Rate (EER) threshold as a reference to compute the error rates. In other words, it adapts a relative difference formula to quantify demographic fairness by comparing individual group performances to a global standard. Authors argue that by incorporating both within-demographic (WDI) and cross-demographic (CDI) interactions, SEDG is able to provides provide better understanding of the magnitude and type of bias making it a versatile measure. For a setup comprising $\mathcal{D}$ set of demographic groups, the SEDG can be obtained as:
@@
—where the subscripts $m$ and $nm$ refer to the distributions of mated and non-mated pairs, respectively.
In [89], Kotwal and Marcel introduced three fairness evaluation measures that emphasize the separation, compactness, and distribution of genuine and impostor scores. Unlike conventional approaches that depend on system accuracy, these measures focus on assessing differential performance without requiring external parameters such as score thresholds. By examining how well the match is, rather than merely determining a match, this approach provides a more nuanced evaluation of demographic fairness. The Separation Fairness Index (SFI) measures the consistency in separation between the mean genuine and impostor scores across groups—higher consistency implies better fairness. The Compactness Fairness Index (CFI) assesses the similarity in intra-group score variability by comparing the standard deviations of genuine and impostor scores across demographics. The Distribution Fairness Index (DFI) evaluates how closely the entire score distribution of each group matches a reference distribution (e.g., the population average), using a statistical divergence measure. Their normal variants are summarized in Eq. 7.
@@
Several studies have evaluated bias and fairness in FR systems using the standard deviation of performance metrics calculated across demographic groups. These metrics include FMR, FNMR, and True Match Rate (TMR)– where higher value in standard deviation corresponds to greater demographic disparities [94, 95, 34, 77, 73, 93]. Another significant metric is the Skewed Error Ratio (SER), which specifically focuses on worst-case error ratios, providing insights into the performance imbalance across groups [93, 73, 96]. Recent competitions [97, 98, 99] exploring the use of synthetic data for FR and bias mitigation have adopted a trade-off performance metric: the mean accuracy adjusted by the standard deviation. This metric aims to ensure that efforts to mitigate bias do not come at the expense of recognition performance. This metric emphasizes the development of FR models that achieve both high recognition performance and fairness across constituent demographic groups.
Table IV provides a brief summary of the fairness assessment metrics discussed in this section. While several metrics listed here capture similar goals: quantifying inter-group performance disparities, they differ in sensitivity, interpretability, and need for threshold selection. Some, that employ differential FMR/FNMR, are easier to relate to operational performance but may not generalize across systems with differing threshold settings. Others, like d-prime and fairness index measures, focus on distributional shifts and are less sensitive to specific thresholds. As accuracy differentials become narrower with improvements in model architectures and training data, future fairness evaluations should include both threshold-based and threshold-agnostic metrics.
V Bias Mitigation in Face Recognition Systems
@@
Pre-processing methods aim to address fairness at the level of training data by modifying or augmenting it to reduce discriminatory patterns or imbalances before the data is used for training. This approach is particularly relevant when representational disparities in the data contribute to demographic bias. The in-processing methods focus on modifications to the FR model during training or fine-tuning phase, often by incorporating constraints or objectives that optimize fairness without significantly compromising recognition accuracy. Finally, post-processing techniques involve adjustments to the output of trained models to ensure fairness across demographic groups. These techniques modify the results, such as classification scores or decision thresholds, to achieve equitable performance without altering the underlying model. Fig. 4 summarizes the categorization of bias mitigation methods– that correlates with the corresponding stages of FR pipeline. These methods are generally aimed at minimizing accuracy disparities across demographic groups while maintaining overall recognition performance. However, empirical evidence from most studies indicates that efforts to reduce such disparities often lead to a reduction in overall accuracy.
V-A Pre-Processing Methods
@@
In [23], Klare et al. showed that FR algorithms were performing worse for female, Black, and younger individuals. To address this concern, they proposed two mitigation strategies based on selection of training data. First, training models on specific demographic cohorts to enhance recognition for those groups, and second, implementing a dynamic face matcher selection approach, where different algorithms, each trained on distinct demographic groups, would be chosen based on probe information. A similar approach was followed by Deb et al. in the context of longitudinal study of FR [102]. Using the data of more than 900 children captured over time, they demonstrated that recognition accuracy decreases as the time gap between image captures grows. To address this, they fine-tuned FaceNet [103] on a separate child face dataset, showing improved accuracy. The authors advocated the need for tailored training and evaluation of FR systems for different age groups. Lu et al. [34] investigated the influence of covariates–such as age, gender, pose, and skin tone– toward the performance of face verification. Their findings showed that using gender information to curate training data improves performance, particularly at low FMRs. In [104], Kortylewski et al. demonstrated the effectiveness of using synthetic face images to mitigate the accuracy differentials arising from real-world datasets. By utilizing synthetic data for pre-training the FR model, they showed that the negative effects of dataset bias, particularly with regard to pose variations, could be significantly reduced. Their experiments revealed that pre-training the CNNs with synthetic data, the need for real-world data can be reduced by up to 75%. Although this study did not directly address demographic fairness, we include it in the review as it highlights the potential of synthetic data in improving the generalization performance of FR systems and reducing (non-demographic) dataset bias. Yucer et al. [105] proposed a GAN-based adversarial image-to-image augmentation method to generate racially balanced training data at subject level while preserving identity features. Evaluated across multiple loss functions, their approach reduced performance variance across racial groups and improved recognition accuracy for minority subjects.
@@
Wang et al. explored the issue of highly-skewed class distributions in FR datasets [106]. They proposed Large Margin Feature Augmentation (LMFA) and Transferable Domain Normalization (TDN) as methods to balance class distributions by augmenting and normalizing the feature space. These methods were shown to enhance the performance of underlying FR models by mitigating issues arising from class imbalance, which often correlates with demographic bias in under-represented groups. In [107], Yin et al. introduced a center-based feature transfer framework to address the under-representation of certain demographic groups in FR datasets. By transferring feature distributions from well-represented classes to under-represented ones, they augmented the feature space for these groups, reducing bias and improving recognition performance for under-represented subjects. Recently, Kotwal and Marcel proposed an Image-to-Image transformation module called Demographic Fairness Transformer (DeFT), which enhances image representations before passing them to pretrained CNNs [108]. The DeFT uses multi-head encoders and soft-attention mechanisms to selectively enhance images based on inferred demographic information. The demographic labels of race or ethnicity are often non-discrete– but this concern has rarely been addressed. In [108], they replaced hard labels with probabilistic weights which are implicitly inferred at run-time. Their experiments show that DeFT reduces bias and improves model fairness, with some models also achieving slightly better accuracy compared to baseline systems.
V-B In-Processing Methods
@@
Gong et al. proposed a group-adaptive classifier (GAC) that uses adaptive convolution kernels and attention mechanisms tailored to different demographic groups [110]. By applying kernel masks and attention maps specific to each group, their method activates facial regions that are more discriminative for each demographic, thereby improving recognition accuracy and fairness across demographic groups. Another approach by the same authors [95] introduced an adversarial network, called DebFace, which utilizes a multi-task learning framework to simultaneously learn identity and demographic attributes. Their method employed adversarial training to disentangle identity features from demographic attributes such as gender, age, and race to effectively reducing differential in the recognition process. Their experiments demonstrated that, for DebFace, not only recognition but also demographic attribute estimation tasks were less biased. In [111], authors proposed unsupervised data partitioning, in iterative manner, followed by Invariant Feature Regularization, to improve generalization across demographic groups. The unsupervised partitioning was chosen as an inexpensive alternative to annotating training data with multiple (confounding) attributes.
@@
Wang et al. [106] 2019 Data-Processing RFW Large-margin feature augmentation to balance class distributions
Yin et al. [107] 2019 Data-Processing LFW, IJB-A, MS-Celeb-1M Feature transfer to enhance under-represented groups
@@
Another approach for domain-specific bias mitigation using disentangled representation learning was proposed by Liang et al. [112]. They introduced a two-stage method combining modules for disentangled representation learning with additive adversarial learning (AAL). While this work does not directly address accuracy differential, it provides useful insights into how domain-specific biases can be mitigated by learning disentangled representations. The effectiveness of this method in reducing bias across various domains suggests its potential applicability in the context of demographic fairness in FR. In [113], a Progressive Cross Transformer (PCT) was proposed to mitigate racial bias by decoupling face representations into identity-related and race-induced components. Using dual cross-transformers, the PCT refines identity features and suppresses racial noise, demonstrating lower racial differentials without compromising recognition accuracy. The concept of score normalization was incorporated as a regularization term into the training objective enabling simultaneous optimization of recognition accuracy and demographic fairness [96]. This was facilitated by constraining the output scores of mated and non-mated pairs to adhere to a pre-defined distribution, and followed by minimizing differences in score distributions across demographic groups. During inference, the overall pipeline did not require modifications as the FR CNN architecture remained unaltered while only the weights were fine-tuned to the new objective.
A couple of works [116, 117] have proposed customized instance margins aimed at preserving consistent error rates– either false positives or false positives combined with true positives. These methods modify the Softmax-based loss by additional penalty terms derived from disparities in the concerned error rates.
@@
V-C Post-Processing Methods
These techniques are applied after the feature extraction and/or matching processes to adjust the final decision scores and reduce bias. We also consider methods that are devised to process features (embeddings) extracted from a pretrained or frozen FR network. While these methods are less commonly used compared to pre- and in-processing techniques, they can still play a role in ensuring equitable outcomes. Additionally, these approaches are relatively easy to integrate into existing systems.

### v3

@@ V Assessment of Demographic Fairness in FR
Since the 2019 Face Recognition Vendor Test (FRVT) [^2] Since 2023, the FRVT initiative has been restructured into the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) programs. report, NIST has included demographic effects in FR algorithms [19]. This assessment involved comparing non-mated pairs within the same demographic group, setting thresholds for algorithms to achieve an FMR of 0.0001 for white males (since this demographic typically associated with the lowest FMR). Constraining non-mated pairs to share the same demographic attributes (e.g., race or gender) is commonly referred to as yoking [92]. The report [19] offered a comprehensive analysis of recognition processes and identified areas where demographic effects might occur. To quantify demographic disparity, NIST initially employed Inequity Ratios (IR), calculating the ratio of maximum to minimum of FMR and FNMR, at score threshold $\tau$, across demographic groups as shown in Eq. 1.
@@
–where, the subscripts $d_{i},d_{j}$ refer to the demographic group; and $\alpha,\beta$ are weighing coefficients. The error rates (FMR and FNMR) have been yoked, meaning they were calculated using pairs exclusively from the same demographic group.
However, considering potentially large range of the error rates, these ratios can become numerically unstable, especially in extreme cases. To alleviate this shortcoming, NIST has considered few modified versions of Inequity Ratios such as adjusting the score threshold ($\tau$), incorporating fixed constants in the denominator, or expressing worst-case error rates relative to arithmetic or geometric means [93]. Using the geometric mean is particularly advantageous due to its extended range over FMR/FNMR values, and this metric, as described in Eq 2 has often been used by NIST.
$\text{A}(\tau)=\frac{\max_{d_{i}}\text{FMR}_{d_{i}}(\tau)}{\text{FMR}^{*}(\tau)}$
$\text{B}(\tau)=\frac{\max_{d_{i}}\text{FNMR}_{d_{i}}(\tau)}{\text{FNMR}^{*}(\tau)},$ (2)
–where the error rates marked with represent the geometric mean computed across all demographic groups. In [94], the authors computed the ratio of each group’s error rate (FMR or FNMR) to the geometric mean across all demographic groups. The overall bias metric was then defined as the sum of the logarithms ($\log 10$) of these ratios for all groups. Another possible approach involves referencing a standard FMR or FNMR in the denominator, which inherently resolves stability issues while providing more robust evaluations of demographic fairness.
It is important to note that the IR metric defined in Eq.1 combines two components: one derived solely from the group-wise FMR values and the other from FNMR values. While the metric multiplies these components after applying user-defined weights ($\alpha$, $\beta$), it is also possible to use the individual terms independently (either or both). The constituent ratios (e.g., the max to min FMR or FNMR across demographic groups, in this case) provide interpretable measures of disparity. The choice of weights typically depends on the specific requirements of the end application. Several prior works have evaluated fairness using single-component metrics focusing on either FMR or FNMR separately[19, 95]. Most metrics, that assess difference in FR performance, discussed below also derive from these two fundamental error types. For brevity, we focus on combined formulations, but note that using one-sided metrics (which is also equivalent to setting one weight to minimum permissible value) is also a valid and reported practice when appropriately motivated.
@@
Mean Absolute Deviation (MAD) - - Evaluates the MAD in specific error rate computed at a fixed threshold.
@@
Schuckers et al. highlighted the importance of accounting for statistical variation when evaluating fairness in FR systems [97]. They noted that the differences among demographic groups can arise either from actual performance disparities or by chance due to sampling variability, leading to potential Type-I errors. To address this, they proposed two statistical methodologies: a bootstrap-based hypothesis test and a simpler test methodology tailored for non-statistical audiences. Their study also conducted simulations to explore the relationship between margin of error and factors such as the number of subjects, attempts, correlation between attempts, underlying FNMRs, and the number of demographic groups. In [94], the authors propose a statistical framework for assessing fairness using Receiver Operating Characteristics (ROC) curves and associated uncertainty measures derived from U-statistics. They argue that the ROC curve is not only central to evaluating system performance but also serves as a key tool for analyzing fairness in similarity scoring. They established asymptotic guarantees for empirical ROC curves and fairness-relevant metrics, emphasizing the importance of accurately quantifying uncertainty. Notably, they highlight that conventional bootstrap methods may misrepresent uncertainty due to the U-statistical nature of FMR and FNMR, and recommend a recentering technique to ensure valid inference. In [99], the researchers at MdTF proposed new metric for assessment of bias based on demographic outcomes. Their metric, GARBE (Gini Aggregation Rate for Biometric Equitability), is inspired by the Gini coefficient– which has a long history of use as a dispersion measure in socio-economic context. The GARBE evaluates statistical dispersion in error rates and emphasizes equitable treatment across demographic groups. Similar to the FDR and IR, this metric combines weighted contributions of FMR and FNMR to produce a single fairness score for a given fixed score threshold. The formula for GABRE and its constituent Gini coefficient are provided in Eq. 4.
@@
Metrics such as GARBE and several others mentioned earlier provide a single scalar value to quantify fairness. This reporting has many advantages as it facilitates simplified, easy comparisons and ranking of systems. However, reducing multiple fairness indicators into a single combined measure as well introduces certain limitations. In particular, the choice of weighting coefficients is often subjective or application-dependent, and may not reflect the actual impact of each component. Moreover, combining ratios (e.g., from FMR and FNMR) that operate on different scales of magnitude can inadvertently skew the metric toward one aspect of fairness, thereby distorting the overall assessment. If the weights are not thoughtfully chosen and justified, such aggregation may obscure significant disparities between demographic groups.
@@
The Sum of Group Error Differences (SEDG) was introduced as the fairness assessment metric address disparities in biometric verification systems in [101]. The SEDG calculates relative deviations in FMR and FNMR across demographic groups from the FMR/ FNMR of global scores. They consider the Equal Error Rate (EER) threshold as a reference to compute the error rates. In other words, it adapts a relative difference formula to quantify demographic fairness by comparing individual group performances to a global standard. Authors argue that by incorporating both within-demographic (WDI) and cross-demographic (CDI) interactions, SEDG is able to provide better understanding of the magnitude and type of bias making it a versatile measure. For a setup comprising $\mathcal{D}$ set of demographic groups, the SEDG can be obtained as:
@@
—where the subscripts $m$ and $nm$ refer to the distributions of mated and non-mated pairs, respectively. The Kolmogorov–Smirnov (KS) bias metric has been adapted for assessing demographic bias in FR by measuring disparities between the score distributions of different demographic groups [103].
In [98], Kotwal and Marcel introduced three fairness evaluation measures that emphasize the separation, compactness, and distribution of genuine and impostor scores. Unlike conventional approaches that depend on system accuracy, these measures focus on assessing differential performance without requiring external parameters such as score thresholds. By examining how well the match is, rather than merely determining a match, this approach provides a more nuanced evaluation of demographic fairness. The Separation Fairness Index (SFI) measures the consistency in separation between the mean genuine and impostor scores across groups–higher consistency implies better fairness. The Compactness Fairness Index (CFI) assesses the similarity in intra-group score variability by comparing the standard deviations of genuine and impostor scores across demographics. The Distribution Fairness Index (DFI) evaluates how closely the entire score distribution of each group matches a reference distribution (e.g., the population average), using a statistical divergence measure. Their normal variants are summarized in Eq. 8.
@@
Several studies have evaluated bias and fairness in FR systems using the standard deviation of performance metrics calculated across demographic groups. These metrics include FMR, FNMR, and True Match Rate (TMR)– where higher value in standard deviation corresponds to greater demographic disparities [104, 105, 41, 86, 83, 100]. The mean absolute deviation (MAD) of a metric across demographic groups is another approach used to quantify bias. For example, in [106], Terhörst et al. evaluated bias by computing the MAD of the TMR at a fixed FMR threshold. Another significant metric is the Skewed Error Ratio (SER), which specifically focuses on worst-case error ratios, providing insights into the performance imbalance across groups [100, 83, 107]. Recent competitions [108, 109, 110] exploring the use of synthetic data for FR and bias mitigation have adopted a trade-off performance metric: the mean accuracy adjusted by the standard deviation. This metric aims to ensure that efforts to mitigate bias do not come at the expense of recognition performance. This metric emphasizes the development of FR models that achieve both high recognition performance and fairness across constituent demographic groups.
Table IV provides a brief summary of the fairness assessment metrics discussed in this section. While several metrics listed here capture similar goals: quantifying inter-group performance disparities, they differ in sensitivity, interpretability, and need for threshold selection. Some, that employ differential FMR/FNMR, are easier to relate to operational performance but may not generalize across systems with differing threshold settings. Others, like d-prime and fairness index measures, focus on distributional shifts and are less sensitive to specific thresholds. An important yet often under-addressed aspect in fairness evaluation is the choice of threshold setting in threshold-based metrics. Most studies adopt a global threshold determined from the overall population FMR or FNMR. Alternatively, some works use yoking, where thresholds are computed using impostor pairs from the same demographic group, aiming to reflect intra-group calibration. In certain cases, thresholds are derived based solely on the FMR of a dominant group (e.g., White Caucasians as seen in FRVT [19]). Some works determine the threshold by fixing the maximum intra-group FMR to enhance robustness against demographic imbalance [95]. Since the threshold setting significantly affects all threshold-based fairness metrics, it is a crucial design choice that must be made with transparency and contextual awareness. As accuracy differentials become narrower with improvements in model architectures and training data, future fairness evaluations should include both threshold-based and threshold-agnostic metrics.
VI Bias Mitigation in Face Recognition Systems
@@
Pre-processing methods aim to address fairness at the level of training data by modifying or augmenting it to reduce discriminatory patterns or imbalances before the data is used for training. This approach is particularly relevant when representational disparities in the data contribute to variations in FR performance across demographic groups. The in-processing methods focus on modifications to the FR model during training or fine-tuning phase, often by incorporating constraints or objectives that optimize fairness without significantly compromising recognition accuracy. Finally, post-processing techniques involve adjustments to the output of trained models to ensure fairness across demographic groups. These techniques modify the results, such as classification scores or decision thresholds, to achieve equitable performance without altering the underlying model. Figure 4 summarizes the categorization of bias mitigation methods– that correlates with the corresponding stages of FR pipeline. These methods are generally aimed at minimizing accuracy disparities across demographic groups while maintaining overall recognition performance. However, empirical evidence from most studies indicates that efforts to reduce such disparities often lead to a reduction in overall accuracy. Moreover, accuracy is often not the most meaningful metric in deployment settings, where systems are evaluated at specific operating points defined by fixing a particular error rate, such as FMR, FNMR, or TMR. As a result, bias mitigation techniques should as well consider application-relevant metrics evaluated at the corresponding operational thresholds.
VI-A Pre-Processing Methods
@@
In [27], Klare et al. showed that FR algorithms were performing worse for female, Black, and younger individuals. To address this concern, they proposed two mitigation strategies based on selection of training data. First, training models on specific demographic cohorts to enhance recognition for those groups, and second, implementing a dynamic face matcher selection approach, where different algorithms, each trained on distinct demographic groups, would be chosen based on probe information. A similar approach was followed by Deb et al. in the context of longitudinal study of FR [113]. Using the data of more than 900 children captured over time, they demonstrated that recognition accuracy decreases as the time gap between image captures grows. To address this, they fine-tuned FaceNet [114] on a separate child face dataset, showing improved accuracy. The authors advocated the need for tailored training and evaluation of FR systems for different age groups. Lu et al. [41] investigated the influence of covariates–such as age, gender, pose, and skin tone– toward the performance of face verification. Their findings showed that using gender information to curate training data improves performance, particularly at low FMRs.
@@
In [115], Kortylewski et al. demonstrated the effectiveness of using synthetic face images to mitigate the accuracy differentials arising from real-world datasets. By utilizing synthetic data for pre-training the FR model, they showed that the negative effects of dataset bias, particularly with regard to pose variations, could be significantly reduced. Their experiments revealed that pre-training the CNNs with synthetic data, the need for real-world data can be reduced by up to 75%. Although this study did not directly address demographic fairness, we include it in the review as it highlights the potential of synthetic data in improving the generalization performance of FR systems and reducing (non-demographic) dataset bias. Yucer et al. [116] proposed a GAN-based adversarial image-to-image augmentation method to generate racially balanced training data at subject level while preserving identity features. Evaluated across multiple loss functions, their approach reduced performance variance across racial groups and improved recognition accuracy for minority subjects. Recently, Kotwal and Marcel proposed an Image-to-Image transformation module called Demographic Fairness Transformer (DeFT), which enhances image representations before passing them to pretrained CNNs [117]. The DeFT uses multi-head encoders and soft-attention mechanisms to selectively enhance images based on inferred demographic information. The demographic labels of race or ethnicity are often non-discrete– but this concern has rarely been addressed. In [117], they replaced hard labels with probabilistic weights which are implicitly inferred at run-time. Their experiments show that DeFT reduces variations in FR performance across demographic groups and, thus improves model fairness, with some models also achieving slightly better accuracy compared to baseline systems.
VI-B In-Processing Methods
@@
Wang et al. explored the issue of highly-skewed class distributions in FR datasets [119]. They proposed Large Margin Feature Augmentation (LMFA) and Transferable Domain Normalization (TDN) as methods to balance class distributions by augmenting and normalizing the feature space. These methods were shown to enhance the performance of underlying FR models by mitigating issues arising from class imbalance, which often correlates with demographic bias in under-represented groups. In [120], Yin et al. introduced a center-based feature transfer framework to address the under-representation of certain demographic groups in FR datasets. By transferring feature distributions from well-represented classes to under-represented ones, they augmented the feature space for these groups, reducing bias and improving recognition performance for under-represented subjects.
Gong et al. proposed a group-adaptive classifier (GAC) that uses adaptive convolution kernels and attention mechanisms tailored to different demographic groups [121]. By applying kernel masks and attention maps specific to each group, their method activates facial regions that are more discriminative for each demographic, thereby improving recognition accuracy and fairness across demographic groups. Another approach by the same authors [105] introduced an adversarial network, called DebFace, which utilizes a multi-task learning framework to simultaneously learn identity and demographic attributes. Their method employed adversarial training to disentangle identity features from demographic attributes such as gender, age, and race to effectively reducing differential in the recognition process. Their experiments demonstrated that, for DebFace, not only recognition but also demographic attribute estimation tasks were less biased. Huang et al. [122] introduced GABN, a de-biasing network that combines gradient attention maps (GAM) with adversarial learning to mitigate racial bias in face recognition. The method enforces consistency of GAMs across racial groups and uses GAM-guided sensitive region erasure to enhance feature learning for darker-skinned individuals. Their experiments showed that GABN was effective at reducing the performance gap between darker-skinned subjects and Caucasians. In [123], authors proposed unsupervised data partitioning, in iterative manner, followed by Invariant Feature Regularization, to improve generalization across demographic groups. The unsupervised partitioning was chosen as an inexpensive alternative to annotating training data with multiple (confounding) attributes.
@@
Alasadi et al. [124] 2019 In-Processing CelebA, UMD Faces Introduces an adversarial learning framework using demographic classifier
@@
Wang et al. [119] 2019 In-Processing RFW Large-margin feature augmentation to balance class distributions
Yin et al. [120] 2019 In-Processing LFW, IJB-A, MS-Celeb-1M Feature transfer to enhance under-represented groups
@@
Georgopoulos et al. [82] 2020 In-Processing KANFace Method to disentangle task-relevant features from bias terms in embeddings
@@
Wang et al. [128] 2021 In-Processing BUPT-Balancedface, Globalface Meta Balanced Network (MBN): a bilevel optimization framework that learns adaptive margin values per skin-tone group
@@
Serna et al. [130] 2022 In-Processing DiveFace, RFW, BUPT-Balancedface A triplet-based discrimination-aware loss, applicable to pre-trained FR models
@@
Huang et al. [122] 2023 In-Processing RFW Enforces consistent gradient attention across demographic groups and follows attention-guided erasure to enhance feature learning
Wang et al. [132] 2023 In-Processing RFW Minimizes performance gaps across identities using MixFair Adapter- without relying on sensitive attribute labels
Dooley et al.  [46] 2023 In-processing VGGFace2, CelebA Uses Neural Architecture Search and Hyperparameter Optimization to discover fair FR architectures
@@
Continued on the next page.
TABLE VI: (Continued)
Reference Year Type of Method Test Dataset Summary
@@
Conti & Clémençon [140] 2024 Post-processing RFW Centroid-based loss aligns subgroup performance to a reference group via a lightweight fairness module
Another approach for domain-specific bias mitigation using disentangled representation learning was proposed by Liang et al. [125]. They introduced a two-stage method combining modules for disentangled representation learning with additive adversarial learning (AAL). While this work does not directly address accuracy differential, it provides useful insights into how domain-specific biases can be mitigated by learning disentangled representations. The effectiveness of this method in reducing bias across various domains suggests its potential applicability in the context of demographic fairness in FR. In [124], the authors introduced an adversarial deep learning framework for face matching aimed at reducing demographic bias. The model incorporated a demographic classifier adversary during training to encourage the learned feature representations to be invariant to the demographic attribute (gender, in their work). Their results showed that the adversarial setup reduced disparities in matching performance across demographic groups while maintaining accuracy comparable to a standard baseline.
In [126], a Progressive Cross Transformer (PCT) was proposed to mitigate racial bias by decoupling face representations into identity-related and race-induced components. Using dual cross-transformers, the PCT refines identity features and suppresses racial noise, demonstrating lower racial differentials without compromising recognition accuracy. In [82], the authors proposed a feature decomposition-based framework that separates task-relevant features from bias-related components in deep embeddings. In addition to FR, they applied this framework to age estimation and gender classification, demonstrating that certain demographic attributes can be disentangled without significantly degrading task performance.
A bias mitigation strategy leveraging Neural Architecture Search (NAS) and Hyperparameter Optimization (HPO) for fair FR models was proposed in [46]. By constructing a tailored search space, their authors synthesized architectures that generalize across datasets and protected attributes, such as gender. The models resulted from this work demonstrated reduced linear separability of demographic attributes.
@@
The concept of score normalization was incorporated as a regularization term into the training objective enabling simultaneous optimization of recognition accuracy and demographic fairness [107]. This was facilitated by constraining the output scores of mated and non-mated pairs to adhere to a pre-defined distribution, and followed by minimizing differences in score distributions across demographic groups. During inference, the overall pipeline did not require modifications as the FR CNN architecture remained unaltered while only the weights were fine-tuned to the new objective.
A couple of works [127, 133] have proposed customized instance margins aimed at preserving consistent error rates– either false positives or false positives combined with true positives. These methods modify the Softmax-based loss by additional penalty terms derived from disparities in the concerned error rates. Wang et al. [128] proposed Meta Balanced Network (MBN), a meta-learning framework that dynamically adjusted margin values for different skin tone groups using gradients from a small, unbiased validation set. By framing margin optimization as an outer loop and model training as an inner loop, MBN learnt adaptive margins that improve fairness across groups. The bilevel optimization enabled their model to reduce bias while maintaining performance. Serna et al. [130] proposed Sensitive Loss, a discrimination-aware extension of the triplet loss function that incorporated a sensitive triplet mining strategy. Designed as an add-on to pre-trained FR models, this method enabled fairness improvements without retraining entire architectures. Wang et al. [132] propose MixFairFace, a framework designed to reduce identity bias– performance inconsistencies across individual identities– rather than relying on demographic group labels. The core component, MixFair Adapter, uses a sample-mixing strategy to estimate and minimize identity-level bias during training. This label-free approach enables fairness enhancement without requiring sensitive attribute annotations like race or gender.
@@
VI-C Post-Processing Methods
These methods are applied after the feature extraction or matching stages to adjust decision scores and mitigate bias. Post-processing approaches generally fall into two categories: those that modify the features (embeddings) and those that operate directly on the similarity scores or their distributions. While these methods are less commonly used compared to pre- and in-processing techniques, they can still play a role in ensuring equitable outcomes. Additionally, post-processing methods are generally easy to integrate into existing FR pipelines. Their advantages include avoiding the need to retrain large models from scratch, enabling fairness adaptation using a smaller dataset, faster training times, and the ability to retain the performance of (a strong) pre-trained backbone. However, since the feature extractor remains fixed, these methods may offer limited capacity to correct upstream bias or improve learned representations.
Morales et al. [84] introduced SensitiveNets, a framework that enforces the removal of sensitive attribute information from learned feature representations. By disentangling sensitive attributes from task-relevant features, their method ensures that downstream agents cannot exploit protected information, thereby enhancing both privacy and fairness. Salvador et al. proposed FairCal: a post-processing method aimed at improving fairness in face verification without retraining or requiring sensitive attribute labels [103]. They clustered image embeddings (obtained from frozen FR network) in an unsupervised manner and applied conditional calibration maps per cluster. They reported reduced false matches while improving overall accuracy. In [95], Conti et al. employed a shallow MLP on top of the frozen FR model as a post-processing strategy to mitigate gender bias in FR. They enhanced the representations of underperforming subgroups using a Fair von Mises-Fisher loss with class-specific variance control. The work in [140] introduced a post-processing method that improves fairness in pretrained (frozen) FR models using a centroid-based loss called Centroid Fairness. A lightweight Fairness Module was trained to align subgroup performance with that of a reference group, without modifying the original model.